there are three consecutive integers such that the sum of the first two integers minus the third integer is four. what are the integers?
An answer provided in text says x=6, however ive calculated x=-2 ... am i missing a step or have i miscalculated something? please help. and thank you in advance. :)

Answer by Maths68(1474)

Let
First integer = x+1
Second integer=x+2
Third integer = x+3

Three consecutive integers such that the sum of the first two integers minus the third integer is four.
(x+1+x+2)-(x+3)=4
2x+3-x-3=4
x=4


First integer = x+1=4+1=5
Second integer=x+2=4+2=6
Third integer = x+3=4+3=7
Three consecutive integers are 5,6,7